OpenStudy (anonymous):

A card is chosen at random from a pack of cards and replaced. This experiment is carried out 520 times. State the expected number of times on which the card is either an ace or a club or both.

OpenStudy (anonymous):

1) P(a club) = 13/52 E( a club) = np = 520*13/52 = 130 2) P(an ace) = 4/52 E(an ace) = 520*4/52 = 40 3) P(a picture card) = 12/52 E(a picture card) = 520*12/52 = 120 4) P(either an ace or a club or both) = 4/52 + 13/52 - 1/52 + 1/52 = 17/52 E(either an ace or a club or both) = 520*17/52 = 170 5) P(neither an ace nor a club) = 1 - 16/52 = 36/52 E(neither an ace nor a club) = 520*36/52 = 360

OpenStudy (anonymous):

The answer is 160... not 170... Not sure how to arrive at that answer though.

OpenStudy (anonymous):

But, 360 is correct

OpenStudy (anonymous):

probability of ace,club or both is 16/52 for 520 draws number times = 16/52 * v 520 = 160

OpenStudy (anonymous):

How did you arrive at that answer?

OpenStudy (anonymous):

well there are 13 clubs and 3 aces (excluding 3 clubs) total 16 on one draw therefore probabily of drawing 1 of these is 16/52 times 520 = 160

OpenStudy (anonymous):

P(club) = 13/52 P(ace) = 4/52 P(club and ace) = 1/52 P(club or ace) = P(club)+P(ace)-P(club and ace) = 16/52

OpenStudy (anonymous):

I thought there were 4 aces?

OpenStudy (anonymous):

yes bu one is a club - so you dont count that twice

OpenStudy (anonymous):

just remember the rule P(A or B) =P(A)+P(B)-P(A and B)

OpenStudy (anonymous):

yep - trust us order!!

OpenStudy (anonymous):

I am. And, it's right what you're writing... it's just you first said there are 3 aces... and I didn't see the exlusion part. Sorry about that. I understand it now :)

OpenStudy (anonymous):

thats ok